[Halld-jlab] Hall D Staff Meeting Notes, October 21, 2013

Mark Ito marki at jlab.org
Mon Oct 21 18:52:30 EDT 2013


Folks,

Find the notes below and at

https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_October_21,_2013

   -- Mark
__________________________________


Private:Hall D Staff Meeting, October 21, 2013

    Present: Vladimir Berdnikov, Eugene Chudakov (chair), Mark Dalton,
    Hovanes Egiyan, Nerses Gevorgyan, Mark Ito, Vanik Kakoyan, Bert
    Manzlak, Lubomir Pentchev, Sergey Pozdniakov, Yi Qiang, Elton Smith,
    Steve Smith, Sascha Somov, Simon Taylor, Ivan Tolstukhin, Elliott
    Wolin, Beni Zihlmann

    Call to order: 3:00 pm

Steve Smith

      * [22]Safety Culture Survey results presented
      * need more detailed responses to move forward with understanding
           + supervisory focus groups will meed
      * low score: lab adapts quickly and learns from errors
      * high score: training is adequate
      * Physics Division has relatively low response rate
      * Elliott: in Physics, questions are not always appropriate
      * Elton: hard to tell where incidents occur when lessons learned are
        circulated

Bert Manzlak

      * incident last week: tech scraped head on uni-strut support while on
        ladder cabling
           + partly situational awareness: conditions in Hall changing
             rapidly
           + work plan had been filed on Hall D List
           + worker responded correctly: went to Medical Services, got
             first aid
           + could have been prevented with bump cap; none in Hall now,
             Bert will bring down a few
      * there will be machine tool assessment
           + site office and EH&S will inspect work locations
           + power saws, lathes, etc. need OSPs, authorization lists
      * Mont and Rolf were in the Hall
           + saw good communication between lead engineer and technical
             staff

Eugene

      * the DNP meeting will proceed as planned
      * for PAC 41, no new proposals, instead they will work on
        prioritization
      * Lab in good standing vis-a-vis recent evaluations
           + NSAC confirmed high priority
           + comparative research review by DOE
                o included universities
                o Lab top in all six categories except for training
                  students (but second in that one)
      * coming up: injector test facility
           + 10 MeV beam, one cryo module, high current
           + detector test capability not well defined
      * accelerator having series of readiness reviews in preparation for
        operations
           + Hall D will be subject to some form of review
           + beam in tagger hall could be there in April
                o at least pulsed beam to dump before ramp of tunnel up to
                  nominal Hall beam level
                o beam to tagger hall depends on power arc power supply
                  delivery; delivery would have to be earlier than expected
           + review may come as early as early January
           + Bert: Bob May is overseeing review process: line of inquiry,
             scope of each review
      * we have to start planning beam commissioning: what instrumentation?
        what are procedures? in what order?
           + Elton is putting together a plan; meetings will be held to
             review plan
      * arrangement from last week limiting payments and charging still in
        effect
           + engineers and designers charging to Ops
                o training on transition from Ideas to NX will be delayed
                  however
      * urged backup of data from magnet tests/mapping
      * money for visitors (Soto, Romanov) not released yet

Beni

      * working on read-out of new F1-TDCs
           + got pulser system running
           + exercises pre-amp to read-out

Elton

      * came back from MENU
           + COMPASS presented a lot of data
           + no results on exotics, perhaps this Fall
           + showed results on new scaler, interferes with a1, explains
             anomalous line shape
           + will be working on proceedings
      * will be working on beam commissioning plan

Hovanes

      * last 2 weeks working on alarms
           + mainly voltage control, thousands of channels to monitor
      * Nerses and Vanik working on BCAL voltage control
           + GUIs exist
      * last week chassis for IU bases came in
           + now Anagate devices mounted in chassis
           + EPICS support needed modification for this configuration
           + testing being done in EEL 126
      * work on beamline instrumentation
           + getting cables, VME crates
           + halo counters from around beamline, in groups of four
           + one group in collimator cave another near target
      * data transfer meeting with Computer Center (CC)
           + Paul Letta installing raid, data transfer scripts
      * meeting with Facilities, CC on fiber installation held
      * computers have arrived
           + 16 nodes in 4 boxes plus one 32-core machine for EMUs
                o some for level 3 monitoring
                o some for event builders
           + installed in racks, not booted yet
           + two 90 TB raids installed (75 TB usable each)
                o infiniband switch installed, need a spar power supply for
                  it
                o Elliott: will have to return borrowed raid we have been
                  using

Sascha

      * meeting with Jarrod Fitzpatrick on diamond contract
           + amended version sent to Richard
           + hopefully award by beginning of Nov.
      * hodoscope counters going well
           + Natalie and Nathan stripping cables
           + Chris and Nick soldering bases
           + test procedure designed
      * pair spectrometer tiles
           + Chuck produced drawings for support plates[?], some
             corrections needed
           + will submit for machining next week
           + electronics boards made, need to be tested, then populated
             with SiPMs
      * trigger
           + meeting with Ben, list of triggers reviews
           + he is working on GTP firmware, should be done in a few weeks
           + Hai working on CTPs, there were some problem boards
           + Elton: what about cosmic triggers for BCAL
                o each module can generate a trigger
                o there are 12 bits available in trigger word after
                  accounting for global energy sum
      * no power in Hall yet
      * controllers for FCAL installed, TIs installed, cables up
      * microscope
           + running out of money
           + first bundle due in a week, must be tested first

Yi

      * Fast DAQ for solenoid
           + PXI module now in office
           + testing, looking into black-out issue, found a bent pin on an
             unused channel
      * DIRC simulation
           + geometry added to HDDS
           + studying background rates, reported at PID upgrade meeting
           + researching different photon detection technologies
      * Vanik and Nerses have demonstrated the BCAL voltage control
           + will write set voltages into database
      * Eugene: plan to change settings on temperature?
           + can be done, software exists
           + want to rely as much as possible on on-board temperature
             compensation using thermister

Mark I.

      * data challenge 2
           + fixed calibration database problem
      * decay chain reporting
           + problem understood; fix on the way
      * working on a new tagged release with a new tag of HDDS
      * N. B. nodes on CUE changing: e. g. jlabl3 changed from RHEL5 to
        RHEL6
      * start counter
           + PO placed for counters and their machining
           + test stand built
           + construction jig design underway
      * TOF
           + modules constructed
           + time resolution looks good
           + shipping crates constructed
           + JLab needs to get the frame built

Sergey

      * FCAL preparations
           + 17 new CPUs, 12 installed in crates, cabled to switch
           + IP address distribution in place
           + method for rebooting implemented
           + two more switches promised by CC
           + two master TIs installed
           + waiting for power, perhaps this week
      * inventory database
           + got new version of spreadsheet from Fernando and uploaded data
           + 1,100 items
           + labeling electronics in progress
           + some will be in F117, some in counting room
           + geographic locations now tracked

Simon

      * working on DC alignment with straight tracks
           + offsets and rotations for each CDC wire
      * rare y decay work
           + plan for workshop in March at UMass

Lubomir

      * FDC will move this week
           + work going on on alignment of rails
           + probably no move before Wed.
           + will flush with nitrogen
      * plan: first do surveys, then trial insertion, more surveys, then
        pull out
      * looking at data from full electronics test
           + need to align strips, measure rotation
           + some groups of wires and groups of strips can be treated
             together
           + tension deforms strips
           + corrections have been computed and applied
           + note that this is only for one package
           + spare cathode will be re-scanned (optical) for geometry
           + will need eventually to calibrate

Mark D.

      * network cabling optimization
           + discussion on plan with CC
      * read out of flashes on test set-up
           + issue with mismatch of firmware with final data format
           + Ed will fix (about a week)
           + will be able to go from boards to CODA data

Elliott

      * data transfer from raid to tape library: conventions need to be
        decided with CC, location, procedure
      * FADC firmware need upgrade
      * new switches needed for hall
      * are there enough ethernet ports in rack room?
           + Paul L. will do inventory
           + could get another blade if necessary
      * Dave Butler moved some computers around (physically and
        network-wise); Beni did some computer moving as well
      * data challenge
           + worked on utilities
           + have a schedule for trigger fibers, Chris Cuevas says mid Nov.
      * DNS server, supplied by CC, we will have access to configuration
      * DHCP server: will be run by us
      * convention for voltage supply IP names decided? not?

    Adjourn 4:30 pm

    Retrieved from [36].

References

   22. 
https://halldweb1.jlab.org/talks/2013-4Q/PHY%20Safety%20Culture%20Data%2010-9-13.pptx
   36. 
https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_October_21,_2013


More information about the Halld-jlab mailing list